Q1.   Obesity is a growing problem in the U.S. and educators are looking to determine the probability of children in their schools becoming obese before they leave elementary school. Suppose you have a population of children entering elementary school in Evansville. The probability of a male child becoming obese before leaving elementary school is 0.4 and the probability of a female child becoming obese before leaving elementary school is 0.35. If you select one male and one female from the population of incoming elementary students, the probability that both the male and the female would become obese before leaving elementary school is 0.14. What is the probability that the male or the female (or both) will become obese before they leave elementary school?
    A.   Pr (male | female)=0.4
    B.   Pr (male ∩ female)=0.14
    C.   It is not possible to calculate with the given information
    D.   Pr (male ∪ female)=0.61

Q 2.   The normal distribution can be described by all of the following, except:
    A.   The area under the curve equals 1.0
    B.   The curve is symmetric around the mean
    C.   It has a bell shape
    D.   The values in the middle are more likely to occur
    E.   Pr (μ-2σ < X < μ+2σ) = 0.99

Q 3.   In a study of esophageal cancer and chewing tobacco, 1000 subjects with cancer and 1000 healthy controls were enrolled. Each subject was asked about their previous use of chewing tobacco. Three hundred and fifty of the cases and 120 of the control had exposure to chewing tobacco. This study design is best described as a:
    A.   Intervention trial
    B.   Cross sectional study
    C.   Case control study
    D.   Case series
    E.   Cohort study
Answer #1

Q1. P(Male) = 0.4

P(female) = 0.35

P(Male n Female ) = 0.14

P(Male or female) = P(Male) + P(Female) +P(Male n Female)

= 0.4 +0.35 -0.14 = 0.61

Answer: D. P(male ∪ female) = 0.61

----------

Q2: Answer: E. P (μ-2σ < X < μ+2σ) = 0.99